Microsoft just paid an annual dividend of $2. Analysts expect that the company will

increase dividends at a rate of 13% per year during the next three years, and then increase

at a constant rate of 4% forever. If the discount rate of Microsoft is 15%, what is the

current price of the stock?
Answer #1

D1=(2*1.13)=2.26

D2=(2.26*1.13)=2.5538

D3=(2.5538*1.13)=2.885794

Value after year 3=(D3*Growth rate)/(Discount rate-Growth rate)

=(2.885794*1.04)/(0.15-0.04)

=27.2838705

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=2.26/1.15+2.5538/1.15^2+2.885794/1.15^3+27.2838705/1.15^3

=$23.73(Approx).